OpenStudy (anonymous):

solve for x and find the measure of each angle. one angle 4x second angle 2x+5 third angle id 3(x-5)

OpenStudy (vishweshshrimali5):

See, the sum of all angles of a triangle is 180 degrees.

OpenStudy (vishweshshrimali5):

What is the sum of given angles in terms of x? That is calculate: 4x + (2x+5) + 3(x-5) = ?

OpenStudy (anonymous):

9x -10 =180 but once put in the place of x they don't come out to 180

OpenStudy (vishweshshrimali5):

They do: see \[\large{9x-10 = 180}\] \[\large{\implies 9x = 190}\] \[\large{\implies x = \cfrac{190}{9}}\] No you can verify this by calculating each angle with the help of above value and calculating their sum , you would get it as 180.

OpenStudy (anonymous):

yeah i did that i got 21 but once i plugged in my x variables i got 189

OpenStudy (vishweshshrimali5):

Let me check

OpenStudy (vishweshshrimali5):

First of all 190/9 is 21.111...

OpenStudy (vishweshshrimali5):

No I checked it , you must get 180

OpenStudy (anonymous):

let me check it again

OpenStudy (anonymous):

and your plugging it in the 21.11 in the equations?